# Robust Offline Imitation Learning from Diverse Auxiliary Data

### v1

@@ Abstract
Offline imitation learning enables learning a policy solely from a set of expert demonstrations, without any environment interaction. To alleviate the issue of distribution shift arising due to the small amount of expert data, recent works incorporate large numbers of auxiliary demonstrations alongside the expert data. However, the performance of these approaches rely on assumptions about the quality and composition of the auxiliary data. However, they are rarely successful when those assumptions do not hold. To address this limitation, we propose Robust Offline Imitation from Diverse Auxiliary Data (ROIDA). ROIDA first identifies high-quality transitions from the entire auxiliary dataset using a learned reward function. These high-reward samples are combined with the expert demonstrations for weighted behavioral cloning. For lower-quality samples, ROIDA applies temporal difference learning to steer the policy towards high-reward states, improving long-term returns. This two-pronged approach enables our framework to effectively leverage both high and low-quality data without any assumptions. Extensive experiments validate that ROIDA achieves robust and consistent performance across multiple auxiliary datasets with diverse ratios of expert and non-expert demonstrations. ROIDA effectively leverages unlabeled auxiliary data, outperforming prior methods reliant on specific data assumptions.
Currently at AWS AI Labs. Work done while the author was at UCR.
@@ 1 Introduction
Integration of deep neural networks in reinforcement learning (RL), coupled with the development of efficient training algorithms, has yielded remarkable performance across a wide variety of sequential decision-making tasks, such as playing games [19, 30, 29] and solving complex robotics tasks [5, 16, 12]. Despite this progress, two challenges still remain: the need for extensive environment interactions [17], and the inherent difficulty in designing reward functions for complex real-world tasks [1].
@@
Imitation learning (IL), where an agent learns directly from task demonstrations, has been employed as one way to tackle the aforementioned challenges [1, 24, 11]. IL methods can be categorized as online or offline. Online IL algorithms rely on experiences gathered from the environment by executing intermediate policies during training [11, 15]. However, online interaction may be infeasible, unsafe, or expensive in many real-world settings. Offline IL provides a safer alternative, where agents learn solely from pre-collected expert demonstrations without environmental interaction. Offline IL methods like behavioral cloning (BC) [3] remove the need for online experience. However, offline imitation remains vulnerable to distribution shifts as a result of error accumulation over time [25].
@@
To address the challenge of distribution shift, recent offline IL methods incorporate a substantial number of auxiliary imperfect demonstrations alongside expert demonstrations. These auxiliary demonstrations are not expected to meet any optimality criteria, encompassing a mix of expert, near-expert, and non-expert trajectories. A recent work, DWBC [37], treats this auxiliary data as a mixture of expert and sub-optimal data, and trains a discriminator for weighted behavioral cloning. On the other hand, DemoDICE [13] performs state-action distribution matching on the auxiliary data as a regularization term, in addition to matching the distribution over the expert set by solving a convex optimization problem in the dual space. Both methods share the assumption that some high-reward behavioral data are present in the auxiliary dataset, and consequently, utilize only these expert transitions for policy learning by filtering out the non-expert trajectories. However, as shown in Fig. 1, their performance fluctuates as the proportion of expert data in the auxiliary set varies, since they fail to leverage the information available in the non-expert data. Although non-expert data may not explicitly provide knowledge of the optimal policy, it contains substantial dynamics information for the agent. In practical scenarios, it is highly unlikely that the quality of the demonstration data in the auxiliary dataset will be known a priori. Thus, we design an offline IL algorithm that is more robust to the demonstration quality in the auxiliary data. Our approach does not make any assumptions about the quality of the auxiliary data and achieves reasonably consistent performance as the proportion of high quality to low quality data varies, as shown in Fig. 1.
@@
Figure 1: Robustness to composition of auxiliary data. Performance of existing offline IL algorithms, such as DWBC [37], varies significantly depending on the amount of high-quality transitions present in the auxiliary data (given expert set is kept fixed). In contrast, ROIDA is more robust, highlighting its ability to extract information even from low-quality transitions. The setup shown here is on the Hopper environment; refer to Sec. 5 for details.
@@
In this paper, we present Robust Offline Imitation from Diverse Auxiliary Data (ROIDA), an algorithm that combines the simplicity of BC with the capability of offline RL to leverage transition data of varying quality in the auxiliary dataset. ROIDA does not impose any assumptions on the composition of the auxiliary dataset, and can effectively utilize diverse auxiliary datasets encompassing different ratios of expert and non-expert demonstrations. To achieve this, we identify potential expert state-action pairs in the auxiliary set and assign large weights to these samples in the subsequent weighted BC objective. This involves two key steps: 1) training a discriminator to distinguish between expert and non-expert transitions using Positive-Unlabeled (PU) learning [6], and 2) applying weighted BC to all state-action pairs in the auxiliary data, with weights derived from importance sampling ratios based on thresholded scores provided by the discriminator. However, as previously mentioned, the auxiliary dataset may lack expert state-action pairs. To address sub-optimal transitions, we perform temporal difference learning using the importance sampling ratios from the discriminator as rewards. This approach aims to guide the policy toward the expert states, thereby improving long-term returns (as measured by the discriminator) on states not observed by experts. This guidance allows ROIDA to extract value from low-quality transitions, in addition to expert behaviors, in contrast to previous works.
@@
Experiments on the D4RL benchmark [7] show that ROIDA consistently achieves high performance across seven environments, using auxiliary datasets with varying proportions of expert data. This consistent success highlights ROIDA’s ability to leverage diverse unlabeled data without assumptions on data quality. In contrast, existing offline IL methods perform well only in selective scenarios that match their specific assumptions about the data composition. Our approach is the first to relax the data quality assumptions of the auxiliary dataset, utilizing it to gain comprehensive knowledge about the expert policy and the environment. No other method fully utilizes both the expert and sub-optimal demonstrations in the auxiliary data for policy learning, leading to suboptimal performance.
@@
$\bullet$ We analyze state-of-the-art (SOTA) offline IL methods that utilize auxiliary data along with a small expert set. Our empirical analysis highlights the unrealistic assumptions of these methods, particularly regarding the composition of the auxiliary set. With this in mind, we design an offline IL algorithm, ROIDA, that addresses the limitations of these different methods and remains robust to the quality of demonstrations in the auxiliary dataset.
$\bullet$ ROIDA incorporates PU learning alongside temporal difference learning to effectively utilize both expert and sub-optimal transitions in the auxiliary data.
$\bullet$ We empirically validate that ROIDA is robust to the quality of the auxiliary data and consistently achieves high performance across different environments.
@@ 2 Related Works
Imitation learning. Imitation learning [27] leverages expert demonstrations to train a policy that successfully mimics the expert’s behavior. A common approach is behavioral cloning (BC) [22, 3], which frames policy learning as a supervised learning problem. However, BC exhibits sub-optimal performance in states distant from the training data [25]. Alternatively, inverse reinforcement learning (IRL) first learns a reward function to explain the demonstrated actions before using it to train a policy through any RL algorithm. While popular IRL algorithms [11, 8, 1, 40] can outperform BC, the majority are online methods requiring a substantial number of environment interactions during training, resulting in poor sample efficiency. To circumvent the need for environment interactions, several offline IRL methods [15, 32, 10] based on adversarial training have been proposed. However, these approaches assume that all demonstrations are equally good, resulting in performance degradation when demonstrations contain sub-optimal data, as in our case.
@@
Learning from noisy demonstrations. Various approaches have been introduced to address the challenge of imitation learning from sub-optimal or noisy experts [36, 33, 4, 34, 26]. However, many of these works rely on strong assumptions about the dataset, such as the expert data dominating the majority of the offline dataset [26] or defining sub-optimality as additive Gaussian noise to the action [33]. Other works assume that trajectories are provided with labels indicating degree of optimality [36, 34] or preference rankings between trajectory pairs [4]. Furthermore, these approaches require environment interactions during learning while we focus on the offline setting.
@@
The offline IL setup with an auxiliary dataset was first explored in DemoDICE [13]. DemoDICE conducts state-action distribution matching over the expert set and introduces a regularization constraint to ensure the learned policy remains close to the behavior policy of the auxiliary dataset. DWBC [37] treats the auxiliary data as a mixture of expert and sub-optimal data, and utilizes positive-unlabeled learning to train a discriminator for weighted behavioral cloning. Both approaches encounter challenges when the auxiliary data is highly sub-optimal and might even exhibit inferior performance compared to counterparts utilizing only expert data. In a recent work [28], the authors propose an offline IL algorithm specifically for cases where no expert data is present in the auxiliary dataset. This approach assigns a reward of 1 to the expert transitions and 0 to all auxiliary transitions, employing an offline RL approach alongside BC. While this design can enhance performance when there is no expert data in the auxiliary dataset, the assignment of zero reward to all auxiliary transitions can lead to poor performance when the proportion of experts in the auxiliary dataset is increased. In practical settings, it is highly unlikely that the data quality in the auxiliary dataset will be known beforehand. With this in mind, we design an offline IL algorithm that addresses the limitations of these different approaches and remains robust to the quality of demonstrations in the auxiliary dataset.
@@
Offline reinforcement learning. Offline RL [17] aims to learn policies by utilizing static offline datasets without requiring additional interactions with the environment. Notably, in offline RL, the training dataset is permitted to contain non-optimal trajectories, and the reward for each state-action-next state transition triplet is known. Our algorithm takes inspiration from a subset of methods within the offline RL literature, specifically those employing filtered advantage weighted regression [21, 35, 20] and behavioral cloning augmented off-policy learning [9] In another recent work, UDS [38] utilizes an auxiliary set without reward labels in addition to the usual reward labeled offline RL dataset. The approach applies zero rewards uniformly to any unlabeled data and can be effective in highly specific offline RL scenarios. Different from this related work, we do not have access to any reward-labeled dataset in the offline IL setting.
@@ 3 Problem Setting
We formulate our problem using the standard fully-observable Markov Decision Process (MDP) framework [31]. An MDP $\mathcal{M}$ is characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma,p_{0})$, where $\mathcal{S}$ denotes the state space, $p_{0}$ denotes the initial state distribution, and $\mathcal{A}$ represents the action space. At each time step $t$, given a state $s_{t}\in\mathcal{S}$, the agent selects an action $a_{t}\in\mathcal{A}$ according to its policy $\pi(a_{t}|s_{t})\in\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ denotes the probability simplex over $\mathcal{A}$. Following the execution of action $a_{t}$, the MDP transitions to a new state $s_{t+1}\in\mathcal{S}$ based on the transition probability $\mathcal{T}(s_{t+1}|s_{t},a_{t})$, while the agent receives a reward $r(s_{t},a_{t})\in\mathbb{R}$. The primary objective for the agent is to maximize the expected discounted reward $\mathbb{E}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t})\right]$ with discount factor $\gamma\in[0,1]$. The state-action distribution of this policy $\pi$ under the transition function $\mathcal{T}$ is defined as $d^{\pi}=(1-\gamma)\sum_{t}\gamma^{t}d^{\pi}_{t}$, where $d^{\pi}_{t}$ is the distribution of $(s_{t},a_{t})$ under $\pi$ at step $t$.
@@ 4.1 Overview
As discussed in Sec. 3, the auxiliary demonstrations are not expected to adhere to any optimality criteria, and can contain a mix of expert, near-expert, and non-expert trajectories. Existing methods rely on certain assumptions about the quality of this data to effectively utilize them. DWBC [37] and DemoDICE [13] assume the presence of high-reward expert data in the auxiliary dataset. In practice, knowing the data quality in the auxiliary dataset beforehand is highly unlikely. Thus, we propose ROIDA to effectively leverage transition data of varying quality in the auxiliary dataset. ROIDA uses two key ideas to achieve this.
@@ 4.2 Learning a reward model
In order to perform both weighted BC and temporal difference learning on the auxiliary dataset, we construct a reward model by training a discriminator $d(s,a)$ to discern between expert and sub-optimal transitions. Unlike prior approaches [11, 13] that treat this as a standard binary classification task, designating all samples from $\mathcal{D}_{O}$ as negative, we opt for PU learning in the discriminator training process. This decision is driven by the potential presence of expert transitions within the auxiliary dataset. As a result, we consider the auxiliary dataset as an unlabeled set, encompassing both positive samples (expert state-action transitions) and varied negative samples (non-expert transitions), with the expert dataset serving as the labeled positive dataset.
The core idea in PU learning is to re-weight the different losses for the positive and the unlabeled data in an effort to derive an estimate of the model loss on negative samples, which is not directly accessible. Due to the limited amount of expert data, we use a non-negative risk estimator described in [14] in order to make the discriminator more robust:
$\min_{d}\ \eta\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log d(s,a)\right]\\ +\max\left(0,\underset{(s,a)\sim\mathcal{D}_{O}}{\mathbb{E}}\left[-\log(1-d(s,a))\right]\right.\\ \left.-\eta\underset{(s,a)\sim\mathcal{D}_{E}}{\mathbb{E}}\left[-\log(1-d(s,a))\right]\right)\ .$ (1)
@@
The specific form of the reward is inspired by DICE (DIstribution Correction Estimation) approaches [13, 18] which try to estimate $\log{\frac{d^{\pi_{E}}(s,a)}{d^{\pi_{O}}(s,a)}}$. This ratio serves as an indicator for the importance of a state-action tuple; higher values mean that the expert often takes the action $a$ at state $s$. While DICE methods perform this estimation by training a simple binary classifier, we use PU learning to train a more robust discriminator to prevent treating expert samples from $\mathcal{D}_{O}$ as negatives.

### v3

@@ Abstract
Offline imitation learning enables learning a policy solely from a set of expert demonstrations, without any environment interaction. To alleviate the issue of distribution shift arising due to the small amount of expert data, recent works incorporate large numbers of auxiliary demonstrations alongside the expert data. However, the performance of these approaches rely on assumptions about the quality and composition of the auxiliary data, and they are rarely successful when those assumptions do not hold. To address this limitation, we propose Robust Offline Imitation from Diverse Auxiliary Data (ROIDA). ROIDA first identifies high-quality transitions from the entire auxiliary dataset using a learned reward function. These high-reward samples are combined with the expert demonstrations for weighted behavioral cloning. For lower-quality samples, ROIDA applies temporal difference learning to steer the policy towards high-reward states, improving long-term returns. This two-pronged approach enables our framework to effectively leverage both high and low-quality data without any assumptions. Extensive experiments validate that ROIDA achieves robust and consistent performance across multiple auxiliary datasets with diverse ratios of expert and non-expert demonstrations. ROIDA effectively leverages unlabeled auxiliary data, outperforming prior methods reliant on specific data assumptions. Our code is available at https://github.com/uditaghosh/roida.
@@ 1 Introduction
Integration of deep neural networks in reinforcement learning (RL), coupled with the development of efficient training algorithms, has yielded remarkable performance across a wide variety of sequential decision-making tasks, such as playing games (Mnih et al., 2015; Silver et al., 2017; 2018) and solving complex robotics tasks (Duan et al., 2016; Levine et al., 2016; Kaufmann et al., 2023). Despite this progress, two challenges still remain: the need for extensive environment interactions (Levine et al., 2020), and the inherent difficulty in designing reward functions for complex real-world tasks (Abbeel & Ng, 2004).
@@
Imitation learning (IL), where an agent learns directly from task demonstrations, has been employed as one way to tackle the aforementioned challenges (Abbeel & Ng, 2004; Ross & Bagnell, 2010; Ho & Ermon, 2016). IL methods can be categorized as online or offline. Online IL algorithms rely on experiences gathered from the environment by executing intermediate policies during training (Ho & Ermon, 2016; Kostrikov et al., 2019). However, online interaction may be infeasible, unsafe, or expensive in many real-world settings. Offline IL provides a safer alternative, where agents learn solely from pre-collected expert demonstrations without environmental interaction. Offline IL methods like behavioral cloning (BC) (Bojarski et al., 2016) remove the need for online experience. However, offline imitation remains vulnerable to distribution shifts as a result of error accumulation over time (Ross et al., 2011).
@@
To address the challenge of distribution shift, recent offline IL methods incorporate a substantial number of auxiliary imperfect demonstrations alongside expert demonstrations. These auxiliary demonstrations are not expected to meet any optimality criteria, encompassing a mix of expert, near-expert, and non-expert trajectories. A recent work, DWBC (Xu et al., 2022), treats this auxiliary data as a mixture of expert and sub-optimal data, and trains a discriminator for weighted behavioral cloning. On the other hand, DemoDICE (Kim et al., 2022) performs state-action distribution matching on the auxiliary data as a regularization term, in addition to matching the distribution over the expert set by solving a convex optimization problem in the dual space. Both methods share the assumption that some high-reward behavioral data are present in the auxiliary dataset, and consequently, utilize only these expert transitions for policy learning by filtering out the non-expert trajectories. However, as shown in Fig. 1, their performance fluctuates as the proportion of expert data in the auxiliary set varies, since they fail to leverage the information available in the non-expert data. Although non-expert data may not explicitly provide knowledge of the optimal policy, it contains substantial dynamics information for the agent. In practical scenarios, it is highly unlikely that the quality of the demonstration data in the auxiliary dataset will be known a priori. Thus, we design an offline IL algorithm that is more robust to the demonstration quality in the auxiliary data. Our approach does not make any assumptions about the quality of the auxiliary data and achieves reasonably consistent performance as the proportion of high quality to low quality data varies, as shown in Fig. 1.
@@
Figure 1: Robustness to composition of auxiliary data. Performance of existing offline IL algorithms, such as DWBC (Xu et al., 2022), varies significantly depending on the amount of high-quality transitions present in the auxiliary data (given expert set is kept fixed). In contrast, ROIDA is more robust, highlighting its ability to extract information even from low-quality transitions. The setup shown here is on the Hopper environment; refer to Sec. 5 for details.
@@
In this paper, we present Robust Offline Imitation from Diverse Auxiliary Data (ROIDA), an algorithm that combines the simplicity of BC with the capability of offline RL to leverage transition data of varying quality in the auxiliary dataset. ROIDA does not impose any assumptions on the composition of the auxiliary dataset, and can effectively utilize diverse auxiliary datasets encompassing different ratios of expert and non-expert demonstrations. To achieve this, we identify potential expert state-action pairs in the auxiliary set and assign large weights to these samples in the subsequent weighted BC objective. This involves two key steps: 1) training a discriminator to distinguish between expert and non-expert transitions using Positive-Unlabeled (PU) learning (Elkan & Noto, 2008), and 2) applying weighted BC to all state-action pairs in the auxiliary data, with weights derived from importance sampling ratios based on thresholded scores provided by the discriminator. However, as previously mentioned, the auxiliary dataset may lack expert state-action pairs. To address sub-optimal transitions, we perform temporal difference learning using the importance sampling ratios from the discriminator as rewards. This approach aims to guide the policy toward the expert states, thereby improving long-term returns (as measured by the discriminator) on states not observed by experts. This guidance allows ROIDA to extract value from low-quality transitions, in addition to expert behaviors, in contrast to previous works.
@@
Experiments on the D4RL benchmark (Fu et al., 2020) show that ROIDA consistently achieves high performance across seven environments, using auxiliary datasets with varying proportions of expert data. This consistent success highlights ROIDA’s ability to leverage diverse unlabeled data without assumptions on data quality. In contrast, existing offline IL methods perform well only in selective scenarios that match their specific assumptions about the data composition. Our approach is the first to relax the data quality assumptions of the auxiliary dataset, utilizing it to gain comprehensive knowledge about the expert policy and the environment. No other method fully utilizes both the expert and sub-optimal demonstrations in the auxiliary data for policy learning, leading to suboptimal performance.
@@
  1. 1.
    We analyze state-of-the-art (SOTA) offline IL methods that utilize auxiliary data along with a small expert set. Our empirical analysis highlights the unrealistic assumptions of these methods, particularly regarding the composition of the auxiliary set. With this in mind, we design an offline IL algorithm, ROIDA, that addresses the limitations of these different methods and remains robust to the quality of demonstrations in the auxiliary dataset.
  2. 2.
    ROIDA incorporates PU learning alongside temporal difference learning to effectively utilize both expert and sub-optimal transitions in the auxiliary data.
  3. 3.
    We empirically validate that ROIDA is robust to the quality of the auxiliary data and consistently achieves high performance across different environments.
@@ 2.1 Imitation learning
Imitation learning (Schaal, 1999) leverages expert demonstrations to train a policy that successfully mimics the expert’s behavior. A common approach is behavioral cloning (BC) (Pomerleau, 1989; Bojarski et al., 2016), which frames policy learning as a supervised learning problem. However, BC exhibits sub-optimal performance in states distant from the training data (Ross et al., 2011). Alternatively, inverse reinforcement learning (IRL) first learns a reward function to explain the demonstrated actions before using it to train a policy through any RL algorithm. While popular IRL algorithms (Ho & Ermon, 2016; Fu et al., 2017; Abbeel & Ng, 2004; Ziebart et al., 2008) can outperform BC, the majority are online methods requiring a substantial number of environment interactions during training, resulting in poor sample efficiency. To circumvent the need for environment interactions, several offline IRL methods (Kostrikov et al., 2019; Swamy et al., 2021; Garg et al., 2021) based on adversarial training have been proposed. However, these approaches assume that all demonstrations are equally good, resulting in performance degradation when demonstrations contain sub-optimal data, as in our case.
2.2 Learning from noisy demonstrations
@@
Various approaches have been introduced to address the challenge of imitation learning from sub-optimal or noisy experts (Wu et al., 2019; Tangkaratt et al., 2020; Brown et al., 2019; Wang et al., 2021; Sasaki & Yamashina, 2021). However, many of these works rely on strong assumptions about the dataset, such as the expert data dominating the majority of the offline dataset (Sasaki & Yamashina, 2021) or defining sub-optimality as additive Gaussian noise to the action (Tangkaratt et al., 2020). Other works assume that trajectories are provided with labels indicating degree of optimality (Wu et al., 2019; Wang et al., 2021) or preference rankings between trajectory pairs (Brown et al., 2019). Furthermore, these approaches require environment interactions during learning while we focus on the offline setting.
@@
The offline IL setup with an auxiliary dataset was first explored in DemoDICE (Kim et al., 2022). DemoDICE conducts state-action distribution matching over the expert set and introduces a regularization constraint to ensure the learned policy remains close to the behavior policy of the auxiliary dataset. DWBC (Xu et al., 2022) treats the auxiliary data as a mixture of expert and sub-optimal data, and utilizes positive-unlabeled learning to train a discriminator for weighted behavioral cloning. Both approaches encounter challenges when the auxiliary data is highly sub-optimal and might even exhibit inferior performance compared to counterparts utilizing only expert data. In a recent work (Shao et al., 2023), the authors propose an offline IL algorithm specifically for cases where no expert data is present in the auxiliary dataset. This approach assigns a reward of 1 to the expert transitions and 0 to all auxiliary transitions, employing an offline RL approach alongside BC. While this design can enhance performance when there is no expert data in the auxiliary dataset, the assignment of zero reward to all auxiliary transitions can lead to poor performance when the proportion of experts in the auxiliary dataset is increased. In practical settings, it is highly unlikely that the data quality in the auxiliary dataset will be known beforehand. With this in mind, we design an offline IL algorithm that addresses the limitations of these different approaches and remains robust to the quality of demonstrations in the auxiliary dataset.
2.3 Offline reinforcement learning
@@
Offline RL (Levine et al., 2020) aims to learn policies by utilizing static offline datasets without requiring additional interactions with the environment. Notably, in offline RL, the training dataset is permitted to contain non-optimal trajectories, and the reward for each state-action-next state transition triplet is known. Our algorithm takes inspiration from a subset of methods within the offline RL literature, specifically those employing filtered advantage weighted regression (Peng et al., 2019; Wang et al., 2020; Nair et al., 2020) and behavioral cloning augmented off-policy learning (Fujimoto & Gu, 2021) In another recent work, UDS (Yu et al., 2022) utilizes an auxiliary set without reward labels in addition to the usual reward labeled offline RL dataset. The approach applies zero rewards uniformly to any unlabeled data and can be effective in highly specific offline RL scenarios. Different from this related work, we do not have access to any reward-labeled dataset in the offline IL setting.
@@ 3 Problem Setting
We formulate our problem using the standard fully-observable Markov Decision Process (MDP) framework (Sutton & Barto, 2018). An MDP $\mathcal{M}$ is characterized by the tuple $(\mathcal{S},\mathcal{A},\mathcal{T},r,\gamma,p_{0})$, where $\mathcal{S}$ denotes the state space, $p_{0}$ denotes the initial state distribution, and $\mathcal{A}$ represents the action space. At each time step $t$, given a state $s_{t}\in\mathcal{S}$, the agent selects an action $a_{t}\in\mathcal{A}$ according to its policy $\pi(a_{t}|s_{t})\in\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ denotes the probability simplex over $\mathcal{A}$. Following the execution of action $a_{t}$, the MDP transitions to a new state $s_{t+1}\in\mathcal{S}$ based on the transition probability $\mathcal{T}(s_{t+1}|s_{t},a_{t})$, while the agent receives a reward $r(s_{t},a_{t})\in\mathbb{R}$. The primary objective for the agent is to maximize the expected discounted reward $\mathbb{E}\left[\sum_{t}\gamma^{t}r(s_{t},a_{t})\right]$ with discount factor $\gamma\in[0,1]$. The state-action distribution of this policy $\pi$ under the transition function $\mathcal{T}$ is defined as $d^{\pi}=(1-\gamma)\sum_{t}\gamma^{t}d^{\pi}_{t}$, where $d^{\pi}_{t}$ is the distribution of $(s_{t},a_{t})$ under $\pi$ at step $t$.
@@ 4.1 Overview
As discussed in Sec. 3, the auxiliary demonstrations are not expected to adhere to any optimality criteria, and can contain a mix of expert, near-expert, and non-expert trajectories. Existing methods rely on certain assumptions about the quality of this data to effectively utilize them. DWBC (Xu et al., 2022) and DemoDICE (Kim et al., 2022) assume the presence of high-reward expert data in the auxiliary dataset. In practice, knowing the data quality in the auxiliary dataset beforehand is highly unlikely. Thus, we propose ROIDA to effectively leverage transition data of varying quality in the auxiliary dataset. ROIDA uses two key ideas to achieve this.
@@ 4.2 Learning a reward model
In order to perform both weighted BC and temporal difference learning on the auxiliary dataset, we construct a reward model by training a discriminator $d(s,a)$ to discern between expert and sub-optimal transitions. Unlike prior approaches (Ho & Ermon, 2016; Kim et al., 2022) that treat this as a standard binary classification task, designating all samples from $\mathcal{D}_{O}$ as negative, we opt for PU learning in the discriminator training process. This decision is driven by the potential presence of expert transitions within the auxiliary dataset. As a result, we consider the auxiliary dataset as an unlabeled set, encompassing both positive samples (expert state-action transitions) and varied negative samples (non-expert transitions), with the expert dataset serving as the labeled positive dataset.
The core idea in PU learning is to re-weight the different losses for the positive and the unlabeled data in an effort to derive an estimate of the model loss on negative samples, which is not directly accessible. Due to the limited amount of expert data, we use a non-negative risk estimator described in (Kiryo et al., 2017) in order to make the discriminator more robust:
$\min_{d}\ \eta\underset{(s,a)\sim{\mathcal{D}}_{E}}{\mathbb{E}}\left[-\log d(s,a)\right]+\max\left(0,\underset{(s,a)\sim\mathcal{D}_{O}}{\mathbb{E}}\left[-\log(1-d(s,a))\right]\right.\left.-\eta\underset{(s,a)\sim\mathcal{D}_{E}}{\mathbb{E}}\left[-\log(1-d(s,a))\right]\right)\ .$ (1)
@@
The specific form of the reward is inspired by DICE (DIstribution Correction Estimation) approaches (Kim et al., 2022; Ma et al., 2022) which try to estimate $\log{\frac{d^{\pi_{E}}(s,a)}{d^{\pi_{O}}(s,a)}}$. This ratio serves as an indicator for the importance of a state-action tuple; higher values mean that the expert often takes the action $a$ at state $s$. While DICE methods perform this estimation by training a simple binary classifier, we use PU learning to train a more robust discriminator to prevent treating expert samples from $\mathcal{D}_{O}$ as negatives.